In the xy plane find the equation of the straight line passing through the point (3,2) with y-intercept 1
Answer by stanbon(75887)

In the xy plane find the equation of the straight line passing through the point (3,2) with y-intercept 1
---
You have two points: (3/2) and (0,1)
slope = (2-1)(3-0) = 1/3
intercept is 1
--------------------
Equation:
y = (1/3)x + 1
===================
Cheers,
Stan H.